OpenStudy (anonymous):

Solve. 2x + 3(x − 2) = 5 ? show work

OpenStudy (anonymous):

2x +3x -6 =5 5x = 11 x =11/5

OpenStudy (anonymous):

First, simplify the problem using the distributive property: \[2x+3x-6=5\]Then, Combine like terms: \[5x-6=5\]Get the x term by itself: \[5x=11\]Divide both sides by 5 to isolate x and get your answer \[x=11/5\] Or, 2.2